Find (f•g)(x) and (g•f)(x)
f(x)= 10x- 5, g(x)= 9- 5x
(f•g)(x)= ?
(g•f)(x)= ?

Answer by ad_alta(240)

1) "Plug" g into f: (f•g)(x)=10(9-5x)-5=85-50x.
2) Similarly, (g•f)(x)=9-5(10x-5)=34-50x.